# 0-supple.tex

### v1

@@ 5.1 Simulations
We first simulate the distributed homogeneous tensor setting described in Section 2.1, where
$\mathcal{T}^{*}=\mathcal{G}\times_{1}U_{1}\times_{2}U_{2}\times_{3}U_{3},\quad U_{j}\in\mathbb{O}^{p_{j}\times r_{j}},\mathcal{G}\in\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}.$
We set the dimensions $p_{1}=p_{2}=p_{3}=p$ and the ranks $r_{1}=r_{2}=r_{3}=r$. The core tensor $\mathcal{G}$ is generated by first sampling a tensor $\widetilde{\mathcal{G}}\in\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}$ with i.i.d. $\mathcal{N}(0,1)$ entries and rescaling it as $\mathcal{G}=\lambda\cdot\widetilde{\mathcal{G}}/\lambda_{\min}\big{(}\widetilde{\mathcal{G}}\big{)}$, where $\lambda_{\min}\big{(}\widetilde{\mathcal{G}}\big{)}$ denotes the minimum singular value of all matricizations $\mathcal{M}_{j}\big{(}\widetilde{\mathcal{G}}\big{)}$. The generation procedure of $\mathcal{G}$ ensures that the minimum singular value of $\mathcal{G}$ is $\lambda$, which is denoted as the signal strength $\lambda_{\min}$ defined in Section 2.3. For $j=1,2,3$, the matrix $U_{j}$ is generated via QR decomposition on a matrix $\widetilde{U}_{j}\in\mathbb{R}^{p_{j}\times r_{j}}$ with i.i.d. $\mathcal{N}(0,1)$ entries. Then we independently generate $\mathcal{Z}_{\ell}$ with i.i.d. $\mathcal{N}(0,\sigma^{2})$ entries and obtain $\mathcal{T}_{\ell}=\mathcal{T}^{*}+\mathcal{Z}_{\ell}$ for $\ell=1,2,\dots,L$.
@@
Figure 2: The estimation errors of different methods under the heterogeneous setting where the tensors share the same core.
@@
Figure 3: The estimation errors of different methods under the heterogeneous setting where the tensors have different cores.
Asymptotic Distribution
Furthermore, we verify the validity of the asymptotic distribution established in Section B in the supplementary material by computing the coverage rates of our proposed distributed Algorithm 4, that is, the rate that the estimated confidence region (17) covers the truth $U_{j}$. As clarified in Section B, we estimate the noise level $\sigma$ by
and estimate the singular value matrix $\Lambda_{j}$ by
(a) $p=50$
(b) $p=100$
Specifically, we choose the confidence level $1-\xi$ to be $0.95$ and report the coverage rates of $\widehat{U}_{1}$ and $\widehat{U}_{\mathrm{pooled},1}$ with $p=50$ and $100$ in Figure 4. For both cases, our proposed estimator performs comparably to the pooled estimator when the SNR $\lambda/\sigma$ is sufficiently large, achieving a high coverage rate around the nominal 95% level. It is worth noting that, compared to Figure 1, the requirement for SNR to achieve the asymptotic normality is more stringent than that to attain the optimal statistical error rate. This observation is consistent with our theoretical results: Corollary 2.1 guarantees the optimal statistical error rate under the condition $\lambda/\sigma\geq\sqrt{prL}$, whereas Theorem B.1, which establishes the asymptotic normality, assumes a stronger condition that $\lambda/\sigma\geq L^{1/2}(pr)^{3/4}$.
@@ 5.2 Real Data Analysis
(a) $\bm{r}=(1,2,2)$
(b) $\bm{r}=(2,4,4)$
Figure 5: Comparison of the reconstruction errors within class 0 of the PROTEINS dataset.
In this section, we illustrate the usefulness of our proposed methods on the PROTEINS dataset (Borgwardt et al., 2005; Morris et al., 2020), which contains graphs representing proteins classified as enzymes or non-enzymes. The dataset consists of 1,113 protein graphs, among which 663 are enzymes encoded as class 0, and 450 are non-enzymes encoded as class 1. Each graph represents the structure of a single protein, where the nodes represent the secondary structure elements (i.e., the helices, sheets, and turns) of the protein, and the edges connect nodes that are neighbors along the amino acid sequence or neighbors in space within the protein structure. Following the procedure in Wen et al. (2024), we employ Topological Data Analysis (TDA) to encode the topological and structural features of each graph into a three-mode tensor of dimensions $2\times 50\times 50$, composed of two $50\times 50$ persistence diagrams constructed by two filtration functions. Since obtaining the ground truth $U_{j}$ for real data is difficult, we evaluate the performance of different methods by the reconstruction error of the estimators. Formally, given a tensor $\mathcal{T}$, the reconstruction error of estimators $\big{\{}\widehat{U}_{j}\big{\}}_{j=1,2,3}$ on $\mathcal{T}$ is defined as
@@
We first apply our proposed distributed algorithm for homogeneous data (Algorithm 1) to the tensors within class 0. Concretely, we randomly select $L$ tensors with label 0 as the training samples, denoted by $\{\mathcal{T}_{\ell}\}_{\ell\in\mathcal{I}}$, where $\mathcal{I}$ is an index set with cardinality $L$. The training samples $\{\mathcal{T}_{\ell}\}_{\ell\in\mathcal{I}}$ are input into Algorithm 1 to obtain estimators $\big{\{}\widehat{U}_{j}\big{\}}_{j=1,2,3}$. Next, we randomly select $L^{\prime}$ tensors other than the training samples as the test set, denoted by $\{\mathcal{T}_{\ell}\}_{\ell\in\mathcal{I}^{\prime}}$ ($\mathcal{I}\cap\mathcal{I}^{\prime}=\emptyset$), and then compute the averaged reconstruction error of $\big{\{}\widehat{U}_{j}\big{\}}$ over the test samples, i.e., $\frac{1}{L^{\prime}}\sum\limits_{\ell\in\mathcal{I}^{\prime}}\mathrm{RE}\big{(}\widehat{U}_{1},\widehat{U}_{2},\widehat{U}_{3};\mathcal{T}_{\ell}\big{)}$. For comparison, we also record the reconstruction errors of two other methods:
@@
Specifically, we choose the ranks $\bm{r}=(r_{1},r_{2},r_{3})$ to be $(1,2,2)$ or $(2,4,4)$, fix $L^{\prime}=100$, and let $L$ range from 1 to 100.
Figure 5 presents the reconstruction errors of the three methods, which are averaged over 200 independent repeats of random sample selection. Note that, for the “single” method, we report the average reconstruction error over all training samples $\mathcal{T}_{\ell}$, that is, $\frac{1}{LL^{\prime}}\sum\limits_{\ell\in\mathcal{I}}\sum\limits_{\ell^{\prime}\in\mathcal{I}^{\prime}}\mathrm{RE}\big{(}\widehat{U}_{1,\ell},\widehat{U}_{2,\ell},\widehat{U}_{3,\ell};\mathcal{T}_{\ell^{\prime}}\big{)}$. As shown in Figure 5, our distributed method significantly reduces the reconstruction error compared to the “single” estimators when the number of tensors in the training set is sufficiently large, which is evident for both $\bm{r}=(1,2,2)$ and $\bm{r}=(2,4,4)$. Moreover, the distributed method achieves comparable performance to the pooled method for $\bm{r}=(2,4,4)$ and outperforms the pooled method when $\bm{r}=(1,2,2)$.
@@ 6 Conclusion
This paper presents innovative distributed tensor PCA methods for both homogeneous and heterogeneous settings, addressing the challenges of analyzing high-dimensional tensor data stored across multiple locations. We develop and theoretically validate algorithms that efficiently aggregate shared low-rank subspaces and identify unique components, enhancing estimation accuracy under various data heterogeneity conditions. Our methods demonstrate significant improvements over traditional approaches in extensive simulations and real-data analyses. Future research will focus on expanding these techniques to more complex distributed frameworks, exploring adaptive algorithms that can dynamically adjust to varying data characteristics and network conditions. Further investigations will also delve into the scalability of these methods, aiming to optimize computational and communication efficiencies in broader distributed systems.
@@
“Distributed Tensor Principal Component Analysis”
@@ Appendix A Notations
Appendix B Statistical Inference in the Distributed Environment
@@
The technical reason behind the two-iteration distributed procedure is to ensure a more precise quantification of the local estimation error. Recall that in Theorem 2.1, we assume the initial estimators $\big{\{}\widehat{U}^{(0)}_{j,\ell}\big{\}}$ have an error rate of the order $O(\sqrt{p}\sigma\lambda_{\min}^{-1})$. However, to establish the asymptotic normality, we need a finer requirement for the initial estimation error, which may not be satisfied by $\big{\{}\widehat{U}_{j,\ell}^{(0)}\big{\}}$ but is satisfied by the local estimators $\{\widehat{U}_{j,\ell}\}$ obtained in Algorithm 1. Therefore, we add one more iteration for each individual tensor in Algorithm 1 to obtain refined local estimators. This two-iteration strategy was first investigated by a recent work Xia et al. (2022) for single tensor decomposition. In Theorem B.1, we further provide the asymptotic distribution for our distributed estimator and show that aggregation in a distributed setting helps improve the asymptotic Mean Squared Error (MSE) from a local rate $O(p^{2}r^{2}\sigma^{4}\lambda_{\min}^{-4})$ to the global optimal rate $O(L^{-2}p^{2}r^{2}\sigma^{4}\lambda_{\min}^{-4})$, when the SNR is sufficiently large.
Concretely, Theorem B.1 shows that, when the SNR satisfies $\lambda_{\min}/\sigma\gtrsim L^{1/2}(pr)^{3/4}$, the squared distance $\rho^{2}(\widehat{U}_{j},U_{j})$ has an asymptotic bias $2\sigma^{2}L^{-1}p_{j}\big{\|}\Lambda_{j}^{-1}\big{\|}_{\rm F}^{2}$ and an asymptotic standard deviation $\sqrt{8p_{j}}\sigma^{2}L^{-1}\big{\|}\Lambda_{j}^{-2}\big{\|}_{\rm F}$. Since $\big{\|}\Lambda_{j}^{-1}\big{\|}_{\rm F}\leq\sqrt{r}\big{\|}\Lambda_{j}^{-1}\big{\|}_{2}\leq\sqrt{r}\lambda_{\min}^{-1}$, the asymptotic MSE of $\rho^{2}(\widehat{U}_{j},U_{j})$ is of the order $O(p^{2}r^{2}\sigma^{4}L^{-2}\lambda_{\min}^{-4})$, which is consistent with the error rate established in Theorem 2.1. Moreover, the pooled estimator $\widehat{U}_{\mathrm{pooled},j}$ (defined in Section 2.1) has the same asymptotic distribution as (16), indicating that our estimator achieves the optimal asymptotic MSE we can obtain in the distributed setting.
Based on Theorem B.1, we can further construct confidence regions for $U_{j}$ using the proposed estimator $\widehat{U}_{j}$. Specifically, given a pre-specified level $1-\xi$, we construct the confidence region for $U_{j}$ as follows,
and the singular value matrix $\Lambda_{j}$ can be estimated by
on any machine $\ell$.
Algorithm 4 Distributed Tensor PCA for Inference

### v3

@@ 5.1 Simulations
We first simulate the distributed homogeneous tensor setting described in Section 2.1, where $\mathcal{T}^{*}=\mathcal{G}\times_{1}U_{1}\times_{2}U_{2}\times_{3}U_{3}$, $U_{j}\in\mathbb{O}^{p_{j}\times r_{j}}$, $\mathcal{G}\in\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}.$ We set the dimensions $p_{1}=p_{2}=p_{3}=p$ and the ranks $r_{1}=r_{2}=r_{3}=r$. The core tensor $\mathcal{G}$ is generated by first sampling a tensor $\widetilde{\mathcal{G}}\in\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}$ with i.i.d. $\mathcal{N}(0,1)$ entries and rescaling it as $\mathcal{G}=\lambda\cdot\widetilde{\mathcal{G}}/\lambda_{\min}\big{(}\widetilde{\mathcal{G}}\big{)}$, where $\lambda_{\min}\big{(}\widetilde{\mathcal{G}}\big{)}$ denotes the minimum singular value of all matricizations $\mathcal{M}_{j}\big{(}\widetilde{\mathcal{G}}\big{)}$. The generation procedure of $\mathcal{G}$ ensures that the minimum singular value of $\mathcal{G}$ is $\lambda$, which is denoted as the signal strength $\lambda_{\min}$ defined in Section 2.3. For $j=1,2,3$, the matrix $U_{j}$ is generated via QR decomposition on a matrix $\widetilde{U}_{j}\in\mathbb{R}^{p_{j}\times r_{j}}$ with i.i.d. $\mathcal{N}(0,1)$ entries. Then we independently generate $\mathcal{Z}_{\ell}$ with i.i.d. $\mathcal{N}(0,\sigma^{2})$ entries and obtain $\mathcal{T}_{\ell}=\mathcal{T}^{*}+\mathcal{Z}_{\ell}$ for $\ell=1,2,\dots,L$.
@@
Figure 2: Estimation errors under heterogeneous settings where tensors share the same core.
@@
Figure 3: Estimation errors under heterogeneous settings where tensors have different cores.
@@ 5.2 Real Data Analysis
(a) PROTEINS
(b) PTC_FM
Figure 4: Comparison of the reconstruction errors within class 0 of the two datasets.
In this section, we illustrate the usefulness of our proposed methods on two real-world molecule datasets, PROTEINS (Borgwardt et al., 2005; Morris et al., 2020) and PRC_FM (Helma et al., 2001; Kriege and Mutzel, 2012). PROTEINS consists of 1,113 protein graphs labeled according to whether they are enzymes. Each graph represents the structure of a single protein, where the vertices represent the secondary structure elements (i.e., the helices, sheets, and turns) of the protein, and the edges connect nodes that are neighbors along the amino acid sequence or neighbors in space within the protein structure. PRC_FM contains 349 compound graphs classified according to the carcinogenicity on female mice, where the vertices represent atoms, and the edges represent chemical bonds. Following the procedure in Wen et al. (2024), Topological Data Analysis (TDA) is employed to encode the topological and structural features of each graph into a three-mode tensor of dimensions $2\times 50\times 50$ for PROTEINS and $5\times 50\times 50$ for PTC_FM, composed of $50\times 50$ persistence images constructed by several filtration functions.
Since obtaining the ground truth $U_{j}$ for real data is difficult, we evaluate the performance of different methods by the reconstruction error of the estimators. Formally, given a tensor $\mathcal{T}$, the reconstruction error of estimators $\big{\{}\widehat{U}_{j}\big{\}}_{j=1,2,3}$ on $\mathcal{T}$ is defined as
@@
Concretely, we randomly select $L$ tensors as the training samples, denoted by $\{\mathcal{T}_{\ell}\}_{\ell\in\mathcal{I}}$, where $\mathcal{I}$ is an index set with cardinality $L$. The training samples $\{\mathcal{T}_{\ell}\}_{\ell\in\mathcal{I}}$ are input into Algorithm 2 to obtain estimators $\big{\{}\widehat{U}_{j}\big{\}}_{j=1,2,3}$. Next, we randomly select $L^{\prime}$ tensors other than the training samples as the test set, denoted by $\{\mathcal{T}_{\ell}\}_{\ell\in\mathcal{I}^{\prime}}$ ($\mathcal{I}\cap\mathcal{I}^{\prime}=\emptyset$), and then compute the averaged reconstruction error of $\big{\{}\widehat{U}_{j}\big{\}}$ over the test samples, i.e., $\frac{1}{L^{\prime}}\sum\limits_{\ell\in\mathcal{I}^{\prime}}\mathrm{RE}\big{(}\widehat{U}_{1},\widehat{U}_{2},\widehat{U}_{3};\mathcal{T}_{\ell}\big{)}$. For comparison, we also record the reconstruction errors of two other methods:
@@
Specifically, we fix $L^{\prime}=100$ and let $L$ range from 1 to 100.
Figure 4 presents the reconstruction errors of the three methods for the two datasets, which are averaged over 200 independent repeats of random sample selection. For the “single” method, we report the average reconstruction error over all training samples $\mathcal{T}_{\ell}$, that is, $\frac{1}{LL^{\prime}}\sum\limits_{\ell\in\mathcal{I}}\sum\limits_{\ell^{\prime}\in\mathcal{I}^{\prime}}\mathrm{RE}\big{(}\widehat{U}_{1,\ell},\widehat{U}_{2,\ell},\widehat{U}_{3,\ell};\mathcal{T}_{\ell^{\prime}}\big{)}$. As shown in Figure 4, our distributed method significantly reduces the reconstruction error compared to the “single” and “pooled” estimators, highlighting the advantage of our approach in correctly integrating different tensors in the presence of heterogeneity. This superior performance is attributed to the fundamental differences in how each method handles tensor heterogeneity. The “pooled” estimator performs direct averaging of the raw tensors, which can be inappropriate given the heterogeneity of core tensors across different tensors, as evidenced by the simulation results in Figure 3. In contrast, our distributed estimator identifies and aggregates the common principal components shared across tensors, which leads to better performance than direct tensor averaging. Notably, this finding suggests that even in scenarios where tensor pooling is technically feasible, a distributed estimator may still be preferred as it more appropriately extracts and integrates the common information. Additional real data experiments on transfer learning tasks are explored in Section F of the supplementary material.
@@ 6 Conclusion
This paper presents innovative distributed tensor PCA methods for both homogeneous and heterogeneous settings, addressing the challenges of analyzing high-dimensional tensor data stored across multiple locations. We develop and theoretically validate algorithms that efficiently aggregate shared low-rank subspaces and identify unique components, enhancing estimation accuracy under various data heterogeneity conditions. Our methods demonstrate significant improvements over traditional approaches in extensive numerical analyses. Future research will focus on expanding these techniques to more complex distributed frameworks, exploring adaptive algorithms that can dynamically adjust to varying data characteristics and network conditions.
Acknowledgment
The authors gratefully thank the editor, the associate editor, and the anonymous referees for their helpful comments and suggestions.
Disclosure Statement
The authors report there are no competing interests to declare.
Funding
Elynn Chen’s research was supported in part by the National Science Foundation under Award ID 2412577. Xi Chen would like to thank the support from National Science Foundation via the grant IIS-1845444.
@@ Acknowledgment
“Distributed Tensor Principal Component Analysis with Data Heterogeneity”
@@ Appendix A Notations
Appendix B Discussions on the Second-Order Term
In this section, we provide an example to show that the second-order term $pr\sigma^{2}\lambda_{\min}^{-2}$ in Theorem 2.1 is not improvable, which implies the necessity of the condition $\sqrt{prL}=o(\lambda_{\min}/\sigma)$ in Corollary 2.1 to achieve the optimal rate $\sqrt{pr/L}\sigma\lambda_{\min}^{-1}$. For simplicity, we consider the case $J=3$, $p_{1}=p_{2}=p_{3}=p$, and $r_{1}=r_{2}=r_{3}=1$. We first define $\mathcal{F}_{p,1,\lambda_{\min}}:=\left\{\mathcal{T}:\mathcal{T}=\lambda_{\min}\times_{1}U_{1}\times_{2}U_{2}\times_{3}U_{3},U_{j}\in\mathbb{O}^{p\times 1},j=1,2,3\right\}$ to be the class of target tensors $\mathcal{T}^{*}$ satisfying model (1). Moreover, define the estimator class $\mathfrak{U}_{C_{1},c_{1},C_{2}}:=\Big{\{}\big{(}\widehat{U}_{1},\widehat{U}_{2},\widehat{U}_{3}\big{)}:\mathbb{P}\big{[}\big{\|}\widehat{U}_{j}\widehat{U}_{j}^{\top}-U_{j}U_{j}^{\top}\big{\|}_{2}\leq C_{2}\sqrt{p}\sigma\lambda_{\min}^{-1}\big{]}\geq 1-C_{1}e^{-c_{1}p},\widehat{U}_{j}\in\mathbb{O}^{p\times 1},j=1,2,3\Big{\}}$, which contains all the possible initial estimators satisfying the assumption in Theorem 2.1 for constants $C_{1}$, $c_{1}$, and $C_{2}$. Then we have the following result.
@@ Appendix B Discussions on the Second-Order Term
As assumed in Theorem 2.1, suppose that $\sqrt{p}\sigma/\lambda_{\min}=o(1)$ and $L=O(p^{c_{3}})$ for some $c_{3}>0$. There exist constants $C_{1}$, $c_{1}$, and $C_{2}>0$, such that, if $\lambda_{\min}/\sigma=o(\sqrt{pL})$, then
$\sup_{\begin{subarray}{c}\mathcal{T}^{*}\in\mathcal{F}_{p,1,\lambda_{\min}}\\ \big{(}\widehat{U}_{1,\ell}^{(0)},\widehat{U}_{2,\ell}^{(0)},\widehat{U}_{3,\ell}^{(0)}\big{)}\in\mathfrak{U}_{C_{1},c_{1},C_{2}}\\ \text{for all }\ell=1,2,\dots,L\end{subarray}}\left\lVert\widehat{U}_{1}\widehat{U}_{1}^{\top}-U_{1}U_{1}^{\top}\right\rVert_{\rm F}\geq\widetilde{c}_{2}p\sigma^{2}\lambda_{\min}^{-2},$ (16)
with probability at least $1-\widetilde{C}_{1}e^{-\widetilde{c}_{1}p}-\widetilde{C}_{1}p^{-L}$ for some constants $\widetilde{c}_{1},\widetilde{C}_{1},\widetilde{c}_{2}>0$, where $\widehat{U}_{1}$ is the estimator for $U_{1}$ obtained from Algorithm 1.
The proof of Theorem B.1 is provided in Section D.1. Recall that Theorem 2.1 establishes that the estimation error $\big{\|}\widehat{U}_{1}\widehat{U}_{1}^{\top}-U_{1}U_{1}^{\top}\big{\|}_{\rm F}$ is bounded above by $O\big{(}\sigma\lambda_{\min}^{-1}(pr)^{1/2}L^{-1/2}+pr\sigma^{2}\lambda_{\min}^{-2}\big{)}$. Under the condition $\sqrt{prL}\lesssim\lambda_{\min}/\sigma$, the first term $\sigma\lambda_{\min}^{-1}(pr)^{1/2}L^{-1/2}$ dominates the second term $pr\sigma^{2}\lambda_{\min}^{-2}$, leading to the minimax optimal error rate in Corollary 2.1. Theorem B.1 further provides a lower bound result: if the condition $\sqrt{prL}\lesssim\lambda_{\min}/\sigma$ is violated, there exists a tensor $\mathcal{T}^{*}$ and a set of initial estimators that satisfy all the assumptions in Theorem 2.1 but result in a sub-optimal rate $pr\sigma^{2}\lambda_{\min}^{-2}$. This result demonstrates that the error bound in Theorem 2.1 is sharp and thus the condition $\sqrt{prL}\lesssim\lambda_{\min}/\sigma$ is necessary for achieving the optimal convergence rate.
Appendix C Further Results for Knowledge Transfer
In this section, we generalize our knowledge transfer methodology and theoretical analysis in Section 4 from the two-machine setting into an arbitrary number of machines $L$. Furthermore, we present results not only for fixed weights but also for data-adaptive weights.
Suppose there are $L$ tensors $\mathcal{T}_{\ell}=\mathcal{T}_{\ell}^{*}+\mathcal{Z}_{\ell}$, where $\ell=1$ denotes the target tensor and $\ell=2,\dots,L$ represent the $L-1$ source tensors. Analogous to Section 4, we assume that
$\mathcal{T}_{\ell}^{*}$ $=\mathcal{G}_{\ell}\times_{1}[U_{1}\;V_{1,\ell}]\times_{2}[U_{2}\;V_{2,\ell}]\cdots\times_{J}[U_{J}\;V_{J,\ell}],\quad U_{j}\in\mathbb{O}^{p_{j}\times r_{j,U}},V_{j,\ell}\in\mathbb{O}^{p_{j}\times r_{j,V,\ell}},\mathcal{G}_{\ell}\in\mathbb{R}^{r_{1,\ell}\times r_{2,\ell}\cdots\times r_{J,\ell}},$
with $U_{j}^{\top}V_{j,\ell}=0$ and $r_{j,\ell}=r_{j,U}+r_{j,V,\ell}$ for $j\in[J]$ and $\ell\in[L]$. Our objective is to estimate the singular subspaces spanned by $U_{j}$ and $V_{j,1}$ of the target tensor $\mathcal{T}_{1}$. The noise tensors $\mathcal{Z}_{\ell}$ are assumed to have i.i.d. $\mathcal{N}(0,\sigma_{\ell}^{2})$ entries for $\ell\in[L]$, thus accommodating different noise levels across the tensors.
Algorithm 4 Transferred Tensor PCA for $L$ Machines
Input: Target tensor $\mathcal{T}_{1}$, source tensors $\mathcal{T}_{2},\dots,\mathcal{T}_{L}$, initial estimators $\big{\{}[\widehat{U}_{1,\ell}^{(0)}\;\widehat{V}^{(0)}_{1,\ell}],\dots,[\widehat{U}_{J,\ell}^{(0)}\;\widehat{V}^{(0)}_{J,\ell}]\big{\}}$, and weights $w_{\ell}$ for $\ell=1,2,\dots,L$ that satisfy $\sum_{\ell}w_{\ell}=1$;
Output: Estimators $\big{\{}\widehat{U}_{1},\widehat{V}_{1,1},\widehat{U}_{2},\widehat{V}_{2,1},\dots,\widehat{U}_{J},\widehat{V}_{J,1}\big{\}}$.